OpenStudy (anonymous):

Is anyone on? :( I have a calculus take home test due tomorrow and several questions. PLEASE reply asap!

OpenStudy (biohazard9064):

let me see one of the questions and I will try to help you best I can

OpenStudy (anonymous):

Thank you! okay let me type

OpenStudy (anonymous):

Use implicit differentiation to find dy/dx. \[Cos(xy) +x^5 = y^5\]

OpenStudy (anonymous):

I know it's a chain rule, but i keep getting stuck near the end

OpenStudy (biohazard9064):

well

OpenStudy (biohazard9064):

Is it multiple choice or blank?

OpenStudy (anonymous):

Blank :( There is one multiple choice though i'll ask later

OpenStudy (anonymous):

\[f'(x)= -Sin(xy).(x.1(y')+y) +5x^4=5y^4y'\] correct?

OpenStudy (biohazard9064):

Im gonna be 100% honest I have no clue sorry

OpenStudy (anonymous):

Aww :(

OpenStudy (anonymous):

Are you good at critical points?

OpenStudy (anonymous):

Absolute extrema's?

OpenStudy (biohazard9064):

no

OpenStudy (biohazard9064):

but I know someone who might be able to help

OpenStudy (anonymous):

That'd be very nice

OpenStudy (biohazard9064):

@mathstudent55 Can you help?

OpenStudy (anonymous):

\[y'=\frac{ Sinx^2y-5x^4 }{ -Sinx^2y-5y^4 } \] Was my final answer. Someone please tell me if I'm correct :(

OpenStudy (anonymous):

good luck cheating punk

OpenStudy (anonymous):

Why so mean? :(

OpenStudy (anonymous):

because u broke con nexus honor code :(

OpenStudy (anonymous):

Can you type the code for me? :(